Use the integrating factor to solve the differential equation. dz/dy = z tan y + sin y
1
Expert's answer
2021-09-07T03:12:13-0400
"\dfrac{dz}{dy}-z\tan y=\sin y"

The integrating factor: "\mu(y)=\cos y"


"\cos y\dfrac{dz}{dy}-z(\cos y)(\tan y)=\cos y\sin y"


"\dfrac{d}{dy}((\cos y)z)=-z\sin y+\cos y\dfrac{dz}{dy}"

Then


"\dfrac{d}{dy}((\cos y)z)+z\sin y-z\sin y=\cos y\sin y"

"d((\cos y)z)=\cos y\sin ydy"

Integrate


"\int d((\cos y)z)=\int\cos y\sin ydy"

"(\cos y)z=\dfrac{1}{2}\sin ^2y+C"

"z=\dfrac{\sin ^2y}{2\cos y}+\dfrac{C}{\cos y}"